You have 7 rock music CDs. 5 indie music CDs and 2 pop music CDs. How many ways can you arrange them? In how many different ways can you arrange them so that the CDs of the same type are continguous? Suppose that you want to give your rock music CDs to your friends. You want to give 3 to max, 2 to John and 2 to Alice. How many ways can you do that?
Answer #1

1)

number of ways to arrange =N(7! ways to arrange rock CDs , 5! ways for indie music CD , 2! for pop music and 3! ways to arrange these 3 types )=7!*5!*2!*3! =5040*120*2*6 =7257600

2)

number of ways to divide CD's =7!/(3!*2!*2! )=210